OpenStudy (anonymous):

Solve: 5(x + 8) = 7(x − 2)

OpenStudy (anonymous):

distribute and get the X on one side

OpenStudy (anonymous):

I distributed both sides. Is that correct?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

so i got 5x+40=7x-14 when i distributed

OpenStudy (anonymous):

Then i would add like terms?

OpenStudy (anonymous):

K. I got what you typed.

OpenStudy (anonymous):

add or subtract depending on what the sign is from each side to get the X on one side.

OpenStudy (anonymous):

12x+40= -14?

OpenStudy (anonymous):

to make it more simple what i did was subtract 5x from 7x.

OpenStudy (anonymous):

since 7x is a positive # you would have to subtract from 5X which makes it -2x

OpenStudy (anonymous):

-2x+ 40= -14 then i would subract 40 to both sides. and get the following: -2x=-54, then would i divide positive 2 or negative 2?

OpenStudy (anonymous):

yes and well its -2x so you would want to divide with -2 so you wouldn't be left with a negative x

OpenStudy (anonymous):

so then i would get x=27.

OpenStudy (anonymous):

yes thats what i got! good job!

OpenStudy (anonymous):

ok thank you for the help.

OpenStudy (anonymous):

no problem!